Find the slope and y-intercept 0f 3x-5y+5=0. Show answers in fractions.
Answer by nerdybill(7384)

Manipulate the equation into the "slope-intercept" form:
y = mx + b
where
m is slope
b is the y-intercept
.
3x-5y+5=0
3x-5y= -5
-5y= -3x -5
5y= 3x + 5
y = (3/5)x + (5/5)
y = (3/5)x + 1
.
By inspection, based on the "slope-intercept" form we have:
slope is 3/5
y-intercept is at (0,1)